This question may make you hungry! :) Please solve and explain so I may use for other problems:
3 burgers and 4 orders of fries total $16.80. 2 burgers and 6 orders of fries total $17.20. How much does one burger cost? How much does one order of fries cost?
Thank you! :)

Answer by checkley75(3666)

3b+4f=1680 multiply this equation by 2 [6b+8f=3360}
2b+6f=1720 multiply this equation by -3 [-6b-18f=-5160] now add:
-----------------------------------
6b+8f=3360
-6b-18f=-5160
---------------------------------
-10f=-1800
f=-1800/-10
f=180 cost of the fries.
3b+4*180=1680
3b+720=1680
3b=1680-720
3b=960
b=960/3
b=320 cost of the burger.
proof
2*320+6*180=1720
640+1080=1720
1720=1720